# FedStrategist: A Meta-Learning Framework for Adaptive and Robust Aggregation in Federated Learning

### v1

@@ Abstract
Federated Learning (FL) offers a paradigm for privacy-preserving collaborative AI, but its decentralized nature creates significant vulnerabilities to model poisoning attacks. While numerous static defenses exist, their effectiveness is highly context-dependent, often failing against adaptive adversaries or in heterogeneous data environments. This paper introduces FedStrategist, a novel meta-learning framework that reframes robust aggregation as a real-time, cost-aware control problem. We design a lightweight contextual bandit agent that dynamically selects the optimal aggregation rule from an arsenal of defenses based on real-time diagnostic metrics. Through comprehensive experiments, we demonstrate that no single static rule is universally optimal. We show that our adaptive agent successfully learns superior policies across diverse scenarios, including a “Krum-favorable” environment and against a sophisticated “stealth” adversary designed to neutralize specific diagnostic signals. Critically, we analyze the paradoxical scenario where a non-robust baseline achieves high but compromised accuracy, and demonstrate that our agent learns a conservative policy to prioritize model integrity. Furthermore, we prove the agent’s policy is controllable via a single “risk tolerance” parameter, allowing practitioners to explicitly manage the trade-off between performance and security. Our work provides a new, practical, and analyzable approach to creating resilient and intelligent decentralized AI systems.
@@ 2.3 The Pillars of Trust in Decentralized AI
Building upon the blockchain foundation, the research community has developed three primary mechanisms, or “pillars”, to construct trustworthy decentralized AI systems:
@@
    Incentive Mechanisms & Tokenomics: Addressing the “why” of participation, these systems use economic principles from game theory and mechanism design to create token-based economies [15, 44, 45, 46, 47, 48]. By rewarding desirable actions and punishing malicious behavior, they aim to make honesty the most profitable strategy [16, 49, 50, 51].
@@
These distinct approaches to trust highlight a fundamental challenge in designing robust FL systems, which we conceptualize as the “Aggregation Trilemma,” illustrated in Fig. 1. A practitioner must often choose between optimizing for performance, robustness to attacks, or robustness to data heterogeneity, as no single static rule excels at all three. This motivates the need for an adaptive framework capable of navigating this trade-off space in real time.
@@ 3 Related Work
Our research on an adaptive, meta-learning controller for federated aggregation is situated at the confluence of three major research streams: 1) robust aggregation mechanisms in federated learning, 2) adaptive and dynamic defense strategies, and 3) the application of reinforcement learning for resource management in distributed systems. This section surveys these domains to precisely situate our contribution within the broader landscape of trustworthy decentralized AI.
3.1 Static Defenses: Robust Aggregation in Federated Learning
@@
The foundational challenge in securing Federated Learning (FL) is defending the global model from malicious updates submitted by Byzantine clients [5, 6]. A significant body of work has focused on developing “robust” aggregation rules that can replace the vulnerable Federated Averaging (FedAvg) [1] algorithm. These static defenses can be broadly grouped into two families based on their underlying philosophy.
The first family relies on robust statistics to identify a central tendency that is resistant to outliers. The most prominent examples are Coordinate-wise Median and Geometric Median [57, 58], which are analyzed theoretically in [59]. These methods operate on the principle that the median of a set of values is significantly less sensitive to extreme outliers than the mean. By computing the median for each parameter of the model updates, these methods can effectively filter the influence of poisoning attacks that submit updates with large, anomalous values. Other statistical approaches, such as Trimmed Mean, which removes a certain percentage of the most extreme updates before averaging, operate on a similar principle [59, 60].
The second family uses geometric heuristics. The most well-known of these is Krum (and its variant, Multi-Krum) [19]. Krum’s core assumption is that the model updates from honest, benign clients will form a tight geometric cluster in the high-dimensional parameter space, while malicious updates will be distant outliers. By calculating all pairwise Euclidean distances and selecting the update that is closest to its $k=n-f-2$ nearest neighbors, Krum aims to identify a “true” update from within the honest cluster.
@@
However, a critical and well-documented weakness of all these static defenses is their predictability and context-insensitivity. Research has shown that their fixed logic can be circumvented. For instance, the “local model poisoning attack” by Fang et al. is specifically designed to craft malicious updates that appear “normal” to distance-based rules like Krum [5]. Furthermore, their effectiveness is highly dependent on the underlying data distribution; Krum’s performance degrades significantly in non-IID settings where the updates from honest clients are naturally far apart, a vulnerability also highlighted in [57, 61, 62, 63]. This inflexibility of static rules, which apply the same logic regardless of the environment, motivates the need for dynamic and adaptive defense strategies.
3.2 Adaptive Defenses and Rule-Switching Mechanisms
Recognizing the limitations of static defenses, a second wave of research has focused on creating adaptive systems. These approaches move beyond a single, fixed aggregation rule and instead attempt to adjust the defense based on observed conditions.
Many systems implement this at the application layer by dynamically filtering clients or re-weighting their contributions. The FoolsGold algorithm, for instance, tracks the history of client contributions and uses cosine similarity to identify and down-weight clients participating in a collusion attack [32]. Other systems use reputation scores, managed by a central server or a blockchain, to exclude free-riding clients or weight their contributions based on past performance [8, 64]. The use of Decentralized Identity (DID) and Verifiable Credentials (VCs) in systems like DID-eFed represents a more advanced form of this client-level adaptation, enabling secure and fine-grained participant selection [43].
More advanced frameworks have begun to explore the concept of dynamic aggregation rule selection [65, 66, 67, 68, 69]. The SAMBA framework is a key example, proposing the use of a multi-armed bandit algorithm to switch between different aggregation rules in response to detected attacks [70]. Similarly, the Attack-Adaptive Aggregation (AAA) system employs a neural network with attention mechanisms to learn an optimal aggregation strategy based on client behaviors [71]. These pioneering systems validate the core premise of our work: that selecting the right defense for the current context is a powerful paradigm.
However, these approaches often face their own challenges. They can be computationally heavyweight (e.g., neural network inference per round), and the stability of systems with complex, non-stationary switching dynamics is a significant theoretical hurdle [72, 73]. Furthermore, their adaptation policies, while dynamic, may still be deterministic and therefore learnable and exploitable by a sufficiently sophisticated second-order adversary, as demonstrated by frameworks like AutoAdapt [20]. Our work builds directly on this foundation but seeks to address these gaps by framing the problem within a more lightweight, explicitly cost-aware framework with a focus on analyzing the resulting policy.
3.3 Reinforcement Learning for Resource Management in FL
@@
The third relevant domain is the application of reinforcement learning (RL) and multi-armed bandits (MAB) to optimize processes in distributed systems. This approach has proven highly effective for resource management problems where an agent must make sequential decisions under uncertainty.
In the context of FL, bandit algorithms have been successfully applied to the client selection problem. Cho et al. introduced “UCB-CS,” a bandit-based strategy that selects clients based on an upper-confidence-bound criterion to accelerate convergence while minimizing extra communication [74]. Wan et al. modeled client selection as a multi-armed bandit to improve robustness against Byzantine attacks in their “MAB-RFL” framework [75]. These works demonstrate the viability of using simple RL agents to optimize aspects of the FL pipeline.
@@
Most relevant to our thesis is the “FedCostAware” framework by Sinha et al., which uses a cost-benefit reward function to schedule FL tasks on cloud infrastructure, explicitly controlling the trade-off between performance and financial expense via a tunable parameter, $\lambda$ [76]. This establishes the principle of using an interpretable hyperparameter to manage economic trade-offs in a learning system.
@@
Our proposed FedStrategist framework creates a novel synthesis of these three research streams. While prior work has used bandits for client selection (MAB-RFL) or cost-awareness for scheduling (FedCostAware), and separate work has explored rule-switching (SAMBA), no existing system has unified these concepts. We are the first to propose the use of a lightweight contextual bandit to explicitly and dynamically manage the cost-versus-robustness trade-off for the specific task of aggregation rule selection, and to provide a deep analysis of the resulting learned policy and its stability. It is at this precise intersection, a practical, tunable, and theoretically grounded controller for federated aggregation, that our work makes its novel contribution.
@@ 4.3 Adversarial Threat Model
    Standard Poisoning Attack: A “loud” attack where the malicious client scales its benign update direction by a large factor (e.g., 5×5$\times$5 ×), creating a high-magnitude update designed to be easily detectable by norm-based defenses but highly disruptive to mean-based aggregators.
@@
    Stealth Poisoning Attack: An “intelligent” attack where the adversary crafts its malicious update to precisely match the average L2 norm of the benign clients from the current round, making it invisible to defenses that rely on detecting magnitude-based outliers.
@@ 4.5 The Commander: A Contextual Bandit for Rule Selection
where the accuracy is measured on a small, server-held proxy validation set, $C_{j}$ is a normalized heuristic cost for the chosen rule $j$, and $\lambda_{cost}$ is the key hyperparameter that defines the system’s “risk tolerance.”
@@ 5.3 Framework Implementation
  • ‘fl_core.py’: Contains the ‘Server’ and ‘Client’ classes. The ‘Server’ class manages the global model, orchestrates rounds, and houses the bandit agent. The ‘Client’ class handles local training. A ‘MaliciousClient’ subclass inherits from ‘Client’ and overrides the training method to inject attacks.
  • ‘aggregation.py’: Implements the “Defense Arsenal” as a set of standalone functions (‘fed_avg’, ‘coordinate_wise_median’, ‘krum’).
  • ‘attacks.py’: Contains the logic for the “standard” and “stealth” poisoning attacks.
  • ‘diagnostics.py’: Implements the functions to compute the state vector metrics.
  • ‘bandit.py’: Contains a standard implementation of the LinUCB algorithm using NumPy and Scikit-learn’s Ridge regression for stable linear modeling.
@@ 6 Results
6.1 Analysis of the Learned Policy
Before presenting the quantitative results of the agent’s performance, it is instructive to visualize the policy it is expected to learn. The agent’s task is to map the 2D state space defined by update similarity and variance to an optimal action. Figure 3 illustrates the conceptual decision regions we hypothesize a successful agent will learn.
Fig 3: Conceptual Visualization of the Learned Policy. This diagram illustrates the idealized decision regions. The agent is expected to learn to select the aggressive ‘FedAvg’ in safe conditions (low variance, high similarity), the geometric ‘Krum‘ against loud attacks (high variance), and the statistical ‘Median‘ when faced with high data heterogeneity (low similarity).
@@ 6.2 Performance under a Standard Poisoning Attack
The initial experiment sought to establish a performance baseline by evaluating the four aggregation strategies (our ‘adaptive‘ agent vs. static ‘FedAvg’, ‘Median‘, and ‘Krum‘) against a standard model poisoning attack. We simulated two scenarios: high data heterogeneity ($\beta=0.1$) and moderate heterogeneity ($\beta=0.5$), both with 5 malicious clients out of 20 total participants.
Fig 4: Model accuracy under a standard poisoning attack with moderate heterogeneity ($\beta=0.5$). The FedStrategist (‘adaptive’) agent learns a policy that surpasses all static baselines.
The results, presented in Fig. 4, demonstrate two key findings. First, no single static rule is universally optimal. ‘Krum’ fails completely in the high-heterogeneity environment, as the natural variance of benign clients makes its distance-based outlier detection ineffective. Second, our ‘adaptive’ agent successfully navigates this trade-off. In both scenarios, it learns to avoid the unstable ‘FedAvg’ and the ineffective ‘Krum’, identifying ‘Median’ as the most reliable static baseline and ultimately achieving superior or comparable performance by learning an optimal switching policy.
@@ 6.3 Robustness to an Adaptive Stealth Adversary
The second experiment was designed to test the resilience of our framework against a more sophisticated adversary. We introduced a “stealth” poisoning attack where malicious clients disguise their updates by normalizing their magnitude to match the average of benign clients, specifically targeting norm-based diagnostic metrics.
Fig 5: Model accuracy under a stealth poisoning attack ($\beta=0.5$). The attack’s regularizing effect creates a complex trade-off between the compromised ‘FedAvg’ and the safer robust methods.
The results, shown in Fig. 5, are highly insightful. The stealth attack has an interesting regularizing effect, causing the non-robust ‘FedAvg’ to achieve the highest raw accuracy, as the per-round damage from the poison is reduced. However, ‘FedAvg’ is still being successfully manipulated. The key result is that our ‘adaptive’ agent, despite one of its primary diagnostic metrics being “blinded,” still outperforms the best static robust aggregator (‘Median’). It successfully learns to rely on its remaining signals (like cosine similarity) to make intelligent decisions, balancing the high speed of ‘FedAvg’ with the safety of ‘Median’. This demonstrates the robustness of our multi-metric state representation and the agent’s ability to adapt to a changing threat landscape.
@@ 6.4 Validation of Generalist Strategy in a Krum-Favorable Environment
Our next experiment was designed to test the agent’s generality and lack of inherent bias. We constructed a scenario theoretically optimal for Krum: low data heterogeneity ($\beta=10.0$) to ensure honest clients are tightly clustered, combined with the “standard” high-magnitude attack to ensure malicious clients are clear geometric outliers.
Fig 6: Model accuracy in a Krum-favorable environment ($\beta=10.0$, standard attack). The adaptive agent correctly identifies and tracks the performance of ‘Krum’ as the optimal static strategy.
As shown in Fig. 6, the results validate our hypothesis perfectly. In this environment, unlike before, Krum is performing significantly better thanks to the favorable setup. But the adaptive agent’s performance demonstrates its intelligence: after an initial exploration phase, its accuracy curve converges to and tracks the optimal Krum baseline almost perfectly. This proves that FedStrategist is a true generalist, capable of learning and adapting to the optimal defensive posture based on the specific context of the federated environment.
@@ 6.5 Controllability of the Agent’s Risk Posture
The final and most important experiment investigates the core claim of our thesis: that the ‘lambda_cost’ parameter in the agent’s reward function, $R_{t}=(\Delta\text{Acc})-(\lambda\times C_{j})$, acts as an interpretable ”risk dial.” We ran the adaptive simulation under the stealth attack scenario with four different values of $\lambda$, representing policies from risk-seeking to risk-averse.
The results are presented in Fig. 7. The top plot shows the accuracy curves, while the bottom plot visualizes the agent’s learned strategy by showing the percentage of rounds each aggregation rule was chosen for a given $\lambda$.
Fig 7: Impact of the $\lambda_{cost}$ parameter on agent performance and strategy under the stealth attack. (Left) Accuracy curves for different $\lambda$ values. (Right) The agent’s chosen rule percentages, demonstrating a clear policy shift from risk-seeking (low $\lambda$) to more conservative (high $\lambda$) behavior.
The findings are definitive. The ‘lambda_cost’ parameter directly and predictably controls the agent’s behavior.
  • Low $\lambda$ (e.g., 0.1): When the cost of defense is low, the agent becomes aggressive and risk-seeking. The bar chart shows it selects the high-speed ‘FedAvg’ over 90% of the time. Its accuracy curve mirrors the ‘FedAvg’ baseline, prioritizing performance over security.
  • High $\lambda$ (e.g., 2.0): When the cost of defense is high, the agent becomes more conservative. While it still favors ‘FedAvg’ in this specific scenario, its selection of ‘Median’ and ‘Krum’ increases significantly. Its accuracy curve detaches from the ‘FedAvg’ baseline and provides a more stable, though slower, learning trajectory.
This experiment confirms that FedStrategist is not a monolithic black box. It is a tunable controller. A practitioner can use the $\lambda$ parameter to explicitly define the system’s risk posture, balancing the need for rapid convergence with the requirement for robust defense against attacks. This provides a practical and powerful tool for deploying federated learning in real-world environments with varying security constraints.

### v2

@@ Abstract
Federated Learning (FL) offers a paradigm for privacy-preserving collaborative AI, but its decentralized nature creates significant vulnerabilities to model poisoning attacks. While numerous static defenses exist, their effectiveness is highly context-dependent, often failing against adaptive adversaries or in heterogeneous data environments. This paper introduces FedStrategist, a novel meta-learning framework that reframes robust aggregation as a real-time, cost-aware control problem. We design a lightweight contextual bandit agent that dynamically selects the optimal aggregation rule from an arsenal of defenses based on real-time diagnostic metrics. Through comprehensive experiments, we demonstrate that no single static rule is universally optimal. We show that our adaptive agent successfully learns superior policies across diverse scenarios, including a ”Krum-favorable” environment and against a sophisticated ”stealth” adversary designed to neutralize specific diagnostic signals. Critically, we analyze the paradoxical scenario where a non-robust baseline achieves high but compromised accuracy, and demonstrate that our agent learns a conservative policy to prioritize model integrity. Furthermore, we prove the agent’s policy is controllable via a single ”risk tolerance” parameter, allowing practitioners to explicitly manage the trade-off between performance and security. Our work provides a new, practical, and analyzable approach to creating resilient and intelligent decentralized AI systems.
@@ 2.3 The Pillars of Trust in Decentralized AI
Building upon the blockchain foundation, the research community has developed three primary mechanisms, or ”pillars”, to construct trustworthy decentralized AI systems:
@@
    Incentive Mechanisms & Tokenomics: Addressing the ”why” of participation, these systems use economic principles from game theory and mechanism design to create token-based economies [15, 44, 45, 46, 47, 48]. By rewarding desirable actions and punishing malicious behavior, they aim to make honesty the most profitable strategy [16, 49, 50, 51].
@@
These distinct approaches to trust highlight a fundamental challenge in designing robust FL systems, which we conceptualize as the ”Aggregation Trilemma,” illustrated in Fig. 1. A practitioner must often choose between optimizing for performance, robustness to attacks, or robustness to data heterogeneity, as no single static rule excels at all three. This motivates the need for an adaptive framework capable of navigating this trade-off space in real time.
@@ 3 Related Work
Our research is situated at the confluence of three major research streams: 1) robust aggregation mechanisms in federated learning, 2) adaptive and dynamic defense strategies, and 3) the application of reinforcement learning for resource management in distributed systems. This section surveys these domains to precisely situate our contribution as a necessary advancement over the current state-of-the-art.
3.1 Static Defenses and Their Inherent Limitations
@@
The foundational challenge in securing Federated Learning (FL) is defending the global model from malicious updates submitted by Byzantine clients [5, 6]. A significant body of work has focused on developing ”robust” aggregation rules to replace the vulnerable Federated Averaging (FedAvg) [1] algorithm. These static defenses can be broadly grouped into two families.
The first relies on robust statistics to identify a central tendency resistant to outliers, such as Coordinate-wise Median and Trimmed Mean [57, 58]. The second uses geometric heuristics, most notably Krum, which selects updates based on their proximity to neighboring updates in the parameter space [19].
@@
However, a critical and well-documented weakness of all static defenses is their predictability and context-insensitivity. Research has shown that their fixed logic can be circumvented by sophisticated adversaries [5]. Furthermore, their effectiveness is highly dependent on the underlying data distribution; for instance, Krum’s performance degrades significantly in non-IID settings where the updates from honest clients are naturally far apart [59]. This inflexibility motivates the need for dynamic and adaptive defense strategies.
3.2 Adaptive Defenses: From Rule-Switching to Contextual Awareness
Recognizing the limitations of static defenses, a second wave of research has focused on creating adaptive systems. The pioneering work in this area is the SARA framework by Hu et al. [60], which was the first to formalize aggregation rule adaptation as a Multi-Armed Bandit (MAB) problem. In SARA, each aggregation rule is an ”arm,” and a UCB algorithm selects the optimal rule based on its historical reward, demonstrating the viability of this approach against dynamic attack patterns.
While foundational, this context-free MAB approach has significant limitations. A standard UCB agent learns the best overall strategy but cannot react to the specific conditions of a single round. It is like a doctor prescribing an antibiotic based on its average success rate over the past year, without considering the patient’s immediate symptoms. This approach can be slow to adapt and is vulnerable to adversaries who can manipulate the environment in ways that make the long-term average reward a poor predictor of short-term success.
Our work, FedStrategist, directly addresses this gap. We advance the state-of-the-art by moving from a simple MAB to a contextual bandit. The core novelty of our framework is the introduction of a lightweight, real-time diagnostic state vector, $S_{t}$. This allows our agent to learn a much richer policy, $\pi(S_{t})\to A_{j}$, that maps the current symptoms of the network to the optimal defensive action. This shift from a context-free to a context-aware agent is a critical step towards creating a more intelligent and responsive defense.
3.3 Reinforcement Learning for Optimization in FL
@@
The third relevant domain is the application of reinforcement learning (RL) and bandits to optimize processes in FL. This approach has proven effective for resource management challenges like client selection, where systems like MAB-RFL and UCB-CS use bandits to select the most promising clients for participation [61, 62].
@@
Most relevant to our thesis is the ”FedCostAware” framework by Sinha et al., which introduces a cost-benefit reward function to manage the financial expense of FL on cloud infrastructure [63]. Their work establishes the principle of using a tunable parameter, $\lambda$, to explicitly manage trade-offs.
@@
Our FedStrategist framework creates a novel synthesis of these research streams. While SARA introduced the MAB for rule selection, its approach is context-free and does not consider the cost of defense. While FedCostAware introduced cost-awareness, it was for client scheduling, not defensive strategy. We are the first to unify these concepts. We propose a lightweight contextual bandit that explicitly and dynamically manages the cost-versus-robustness trade-off for the specific task of aggregation rule selection, and provide a deep analysis of the resulting learned policy and its stability. It is at this precise intersection—a practical, tunable, and context-aware controller for federated aggregation—that our work makes its novel contribution.
@@ 4.3 Adversarial Threat Model
    Standard Poisoning Attack: A ”loud” attack where the malicious client scales its benign update direction by a large factor (e.g., 5×5$\times$5 ×), creating a high-magnitude update designed to be easily detectable by norm-based defenses but highly disruptive to mean-based aggregators.
@@
    Stealth Poisoning Attack: An ”intelligent” attack where the adversary crafts its malicious update to precisely match the average L2 norm of the benign clients from the current round, making it invisible to defenses that rely on detecting magnitude-based outliers.
@@ 4.5 The Commander: A Contextual Bandit for Rule Selection
where the accuracy is measured on a small, server-held proxy validation set, $C_{j}$ is a normalized heuristic cost for the chosen rule $j$, and $\lambda_{cost}$ is the key hyperparameter that defines the system’s ”risk tolerance.”
@@ 5.3 Framework Implementation
  • ’fl_core.py’: Contains the ’Server’ and ’Client’ classes. The ’Server’ class manages the global model, orchestrates rounds, and houses the bandit agent. The ’Client’ class handles local training. A ’MaliciousClient’ subclass inherits from ’Client’ and overrides the training method to inject attacks.
  • ’aggregation.py’: Implements the ”Defense Arsenal” as a set of standalone functions (’fed_avg’, ’coordinate_wise_median’, ’krum’).
  • ’attacks.py’: Contains the logic for the ”standard” and ”stealth” poisoning attacks.
  • ’diagnostics.py’: Implements the functions to compute the state vector metrics.
  • ’bandit.py’: Contains a standard implementation of the LinUCB algorithm using NumPy and Scikit-learn’s Ridge regression for stable linear modeling.
@@ 6 Results
6.1 Conceptualizing the Learned Policy
Before presenting the quantitative results, it is instructive to visualize the policy our agent is expected to learn. The agent’s task is to map the diagnostic state space, primarily defined by update similarity and variance, to an optimal action. Figure 3 illustrates the idealized decision regions we hypothesize a successful agent will learn, forming the basis for our experimental validation.
Fig 3: Conceptual Visualization of the Learned Policy. This diagram illustrates the idealized decision regions. The agent is expected to learn to select the aggressive ’FedAvg’ in safe conditions (low variance, high similarity), the geometric ’Krum’ against loud attacks (high variance), and the statistical ’Median’ when faced with high data heterogeneity (low similarity).
@@ 6.2 Performance under a Standard Poisoning Attack
Our initial experiment established a performance baseline against a standard, high-magnitude model poisoning attack. We evaluated four strategies—our ’adaptive’ agent versus static ’FedAvg’, ’Median’, and ’Krum’—across high ($\beta=0.1$) and moderate ($\beta=0.5$) data heterogeneity, with 5 malicious clients out of 20.
Fig 4: Model accuracy under a standard poisoning attack. Top: High heterogeneity ($\beta=0.5$). Bottom: Moderate heterogeneity ($\beta=0.5$). The FedStrategist (’adaptive’) agent learns a policy that surpasses all static baselines.
The results, presented in Fig. 4, yield two key findings. First, no single static rule is universally optimal. ’Krum’ fails completely in the high-heterogeneity environment, as its geometric assumptions are violated. Second, our ’adaptive’ agent successfully navigates this trade-off. In both scenarios, it learns to avoid the unstable ’FedAvg’ and the ineffective ’Krum’, identifying ’Median’ as the most reliable static baseline and ultimately achieving superior performance by learning an optimal switching policy.
@@ 6.3 Robustness to an Adaptive Stealth Adversary
The second experiment tested the system’s resilience against a more sophisticated adversary. We introduced a ”stealth” poisoning attack where malicious clients disguise their updates by normalizing their magnitude, specifically targeting norm-based diagnostic metrics.
Fig 5: Model accuracy under a stealth poisoning attack ($\beta=0.5$). The attack’s regularizing effect creates a complex trade-off between the compromised ’FedAvg’ and the safer robust methods.
The results, shown in Fig. 5, are highly insightful. The attack’s regularizing effect caused the non-robust ’FedAvg’ to achieve the highest raw accuracy. However, this model is fundamentally compromised. The key result is that our ’adaptive’ agent, despite one of its primary diagnostic metrics being ”blinded,” still learned a policy that outperformed the best static robust aggregator (’Median’). It successfully learned to rely on its remaining signals (like cosine similarity) to make intelligent decisions, demonstrating the resilience of our multi-metric state representation.
@@ 6.4 Validation of Generalist Strategy in a Krum-Favorable Environment
Our third experiment was designed to test the agent’s generality and lack of inherent bias. We constructed a scenario theoretically optimal for Krum: low data heterogeneity ($\beta=10.0$) and a ”loud” standard attack.
Fig 6: Model accuracy in a Krum-favorable environment ($\beta=10.0$, standard attack). The adaptive agent correctly identifies and tracks the performance of ’Krum’ as the optimal static strategy.
As shown in Fig. 6, the results validate our hypothesis perfectly. In this environment, Krum is the clear champion among static rules. The ’adaptive’ agent’s performance demonstrates its intelligence: after an initial exploration phase, its accuracy curve converges to and tracks the optimal Krum baseline. This proves that FedStrategist is a true generalist, capable of learning the optimal defensive posture based on the specific context of the federated environment.
@@ 6.5 Controllability of the Agent’s Risk Posture
Our final experiment validates the core claim that the $\lambda_{cost}$ parameter in the agent’s reward function, $R_{t}=(\Delta\text{Acc})-(\lambda\times C_{j})$, acts as an interpretable ”risk dial.” We ran the adaptive simulation under the stealth attack scenario with four different values of $\lambda$.
Fig 7: Impact of the $\lambda_{cost}$ parameter on agent performance and strategy under the stealth attack. (Left) Accuracy curves for different $\lambda$ values, demonstrating a clear shift from risk-seeking (tracking ’FedAvg’) to more conservative behavior. (Right) The agent’s chosen rule percentages, confirming the learned policy change.
The findings, presented in Fig. 7, are definitive. The $\lambda_{cost}$ parameter directly and predictably controls the agent’s behavior.
  • Low $\lambda$ (Risk-Seeking): When the cost of defense is negligible, the agent learns an aggressive policy. The bar chart shows it selects the high-speed ’FedAvg’ over 90% of the time, and its accuracy curve mirrors the ’FedAvg’ baseline.
  • High $\lambda$ (Risk-Averse): When the cost of defense is high, the agent learns a more conservative policy. Its accuracy curve detaches from the ’FedAvg’ baseline, providing a more stable, though slower, learning trajectory, and its selection of robust rules like ’Median’ and ’Krum’ increases.
This experiment confirms that FedStrategist is not a monolithic black box but a tunable controller. A practitioner can use the $\lambda$ parameter to explicitly define the system’s risk posture, providing a practical and powerful tool for deploying federated learning in real-world environments with varying security constraints.